OpenStudy (lilai3):

HOOOO <3333333333333s Peeetttssss????? Irene has 6 pets, some dogs and some birds. Her animals have a total of 16 legs. How many of each pet does Irene have?

OpenStudy (ajprincess):

Let the number of dogs be d and let the number of birds be b. d+b=6-(1) 4d+2b=16-(2) (1)*2=>2d+2b=12-(3) (2)-(3)=>2d=4 d=4/2 d=2 (1)=>2+b=6 b=6-2 b=4 therefore number of dogs is 2 and number of birds is 4.

OpenStudy (lilai3):

thank you, @ajprincess! and i like ur picture of cinderella, by the way. Uhm...thank you!

OpenStudy (ajprincess):

U r welcome @lilai3. thanx for bcoming my fan and also for likng my pic.

OpenStudy (lilai3):

if this wuz facebook, i would've typed (y)....a thumbs up symbol thanks again...and np to ur...stuffz :DDD <333

OpenStudy (ajprincess):

:D

OpenStudy (lilai3):

lol i have a feeling this is nevah gonna end.... :)

OpenStudy (lilai3):

okay...it looks like it ended..lol :DDDD

OpenStudy (ajprincess):

4d+ (12-2d) =16 4d+12-2d=16 2d+12=16 2d=16-12 2d=4 d=4/2 d=2

OpenStudy (ajprincess):

 Nw is this clear?